# PersonaFlow: Boosting Research Ideation with LLM-Simulated Expert Personas

### v1

@@ 8. Conclusion
In this study, we introduced PersonaFlow, an LLM-based research ideation system that aims to address the gap of a lack of access to domain experts during interdisciplinary research ideation. The results showed that working with diverse personas significantly heightened perceived RQ and critique quality without raising the cognitive load. Participants adapted personas across perspectives and the ability to edit persona characteristics increased participants’ feeling of agency, improving the recall of concepts from unknown domains. Participants employed the system for different use cases, showing different cognitive approaches in formulating research ideas, which highlights the versatility and effectiveness of PersonaFlowas a tool for promoting research ideation. Moreover, the persona editing feature allowed for deeper insights into the users’ customization behaviors and what those mean for leveraging AI in research ideation. The results thus provide rich insight into both the opportunities and challenges of using LLM-simulated personas for research ideation, and hence point to future directions for improving human-AI collaboration in interdisciplinary research ideation.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. 2119589. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Additionally, results presented in this paper were obtained using CloudBank (Norman et al., 2021), which is supported by the National Science Foundation under award No. 1925001.
@@ A.1. Participants Demographics
P1 HCI Doctoral 3-4 years
P2 HCI Doctoral 3-4 years
P3 ML/NLP Doctoral 3-4 years
P4 HCI Doctoral 3-4 years
P5 HCI Doctoral 3-4 years
P6 HCI Doctoral (Post) 5+ years
@@
P15 Biomedical Research Doctoral (Post) 5+ years
P16 HCI Doctoral 3-4 years
@@
P18 Economics Master 1-2 years
P19 HCI Doctoral 3-4 years
P20 Bioinformatics Doctoral 3-4 years
P21 HCI Master 1-2 years
@@ A.3. Post-Session Survey Questions (5-point Likert Scale)
The generated literature by personas is domain-relevant.
@@
A.4. Prompts Used in the System

### v2

@@ 8. Conclusion
In this study, we introduced PersonaFlow, an LLM-based research ideation system that aims to address the gap of a lack of access to domain experts during interdisciplinary research ideation. The results showed that working with diverse personas significantly improved perceived relevance and creativity of RQ and critique without raising users’ cognitive load. We also found that PersonaFlow promoted participants’ critical thinking activities under different use cases, showing different cognitive approaches in formulating research ideas. Participants adapted personas across perspectives and the ability to edit persona characteristics increased participants’ feeling of agency, improving the recall of concepts from unknown domains. Moreover, the persona editing feature allowed for deeper insights into the users’ customization behaviors and what those mean for leveraging AI in research ideation. The results thus provide insights into both the opportunities and challenges of using LLM-simulated personas for research exploration, and hence provide design implications and point to future directions for improving human-AI collaboration in interdisciplinary research ideation and beyond.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant #2119589. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. The authors would also like to thank the anonymous reviewers for their valuable feedback. Additionally, results presented in this paper were obtained using CloudBank (Norman et al., 2021), which is supported by the National Science Foundation under award #1925001.
@@ A.1. Participants Demographics
T1 HCI Doctoral 3-4 years
T2 HCI Doctoral 3-4 years
T3 ML/NLP Doctoral 3-4 years
T4 HCI Doctoral 3-4 years
T5 HCI Doctoral 3-4 years
T6 HCI Doctoral (Post) 5+ years
@@
P25 Biomedical Research Doctoral (Post) 5+ years
P22 HCI Doctoral 3-4 years
@@
P20 Economics Master 1-2 years
P29 HCI Doctoral 3-4 years
P30 Bioinformatics Doctoral 3-4 years
P31 HCI Master 1-2 years
@@ A.3. Post-Session Survey Questions (5-point Likert Scale)
The generated literatures by personas are domain relevant.
@@
A.4. Sample Workflow Analyzed During Case Studies
Figure 10. A larger version of the workflow visualization of P20’s thought processes.
A.5. Prompts Used in the System